OpenStudy (anonymous):

What are the zeroes of y = x2 – 2x – 15? –3, –5 3, 5 –3, 5 3, –5

OpenStudy (anonymous):

y = x^2 – 2x – 15 we need to factor this equation or use the quadratic equation. can you do either one of these?

OpenStudy (anonymous):

(x - 3) (x + 5) (-3, 5)?

OpenStudy (anonymous):

No, because you need to make the whole side of the equation equal to zero. What would make (x - 3) zero?

OpenStudy (anonymous):

and also, you factored wrong. that would be correct if it was y = x2 + 2x – 15, but it's a negative 2x, your close, try again.

OpenStudy (anonymous):

"(x - 3) (x + 5) " "(-3, 5)?" close!! but if we FOILed this out what do we get? (x - 3) (x + 5) x^2 +5x -3x -15 x^2 +2x -15 the sign on the 2x is positive, and we need a negative, so if we did (x + 3) (x - 5) x^2 -5x +3x -15 x^2 -2x -15 so, the zeros are 0 = (x+3) 0 = (x-5) x = -3 and 5 so, you have the right idea, just got a little mixed up. ^_^

OpenStudy (anonymous):

*sorry, open study isn't working at all today for me!!